OpenStudy (anonymous):

HELP ME ILL GIVE MEDALS!!!!! What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6? y=-3/2x+8 y=3/2x-4 y=1/2x y=2/3x-4

OpenStudy (anonymous):

help help pleasee!!!!!

OpenStudy (mathmale):

Principle: two lines that are parallel have the same slope. Task: Find the slope of the given line (3x – 2y = –6). Write the point-slope formula for the equation of a straight line. We'll use that formula to find the equation of the new, parallel line.

OpenStudy (anonymous):

sooo what the answwr

OpenStudy (mathmale):

I'd be glad to guide you towards finding the solution of this problem, but it's against the rules for you to ask just for "the answer" and against the rules for me to just provide you with "the answer." Let me know if you're wiling to work through this problem or not.

OpenStudy (***[isuru]***):

first convert the function u got in to the general format of a straight line y = mx + b where m = slope b = y-intercept u got the function 3x - 2y = -6 and u need a line parallel to it and passing through the points ( 4 , 2) first convert 3x - 2y = -6 in to y = mx + b format.. when 2 lines are parallel there slopes are equal... so the slope of the line u want also will be the slope of this line... ( lets say it m) then take the given coordinates ( 4 , 2) u can use the point - slope format here if u take x1 = 4 y1 = 2 then the line u want is \[m ( x - x_{1} ) = y - y_{1}\]

OpenStudy (anonymous):

so the answer is a ?

OpenStudy (anonymous):

yes i am

OpenStudy (***[isuru]***):

yep :)

OpenStudy (mathmale):

Ok, then. Isuru has done a very nice job of explaining the method involved here. You are given 3x - 2y = -6 and are asked to re-write it in the format y=mx+b. I'd bet y ou've done this before. Care to give it a try? Don't be afraid to make mistakes; just get started and show us the work you've done.

OpenStudy (mathmale):

(from ISURU;) first convert 3x - 2y = -6 in to y = mx + b format.. Hint: Solve the first equation for y.